OpenStudy (anonymous):

Which polynomial identity will prove that 65 = 81 - 16? Binomial Multiplication Difference of Cubes Sum of Cubes Difference of Squares

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

81 = 9^2 hint2 : 16 = 4^2 65 = 81-16 = 8^2 - 4^2 = (9+4)(9-4) = (13)(6) = 65

OpenStudy (akashdeepdeb):

It's not B. What do you think the answer should be and why? Try to reason, logically.

OpenStudy (anonymous):

I have no idea because none of those numbers are squared or cubed but theyre being subtracted so I would go with the difference of. So B or D

OpenStudy (akashdeepdeb):

Yes, the answer is D. This is because 81 and 16 are perfect squares. That means, \(9^2 = 81\) \(4^2 = 16\) \(65 = 81 - 16 = 9^2 - 4^2 = difference ~~of~~ perfect ~~squares = D\)